OpenStudy (anonymous):

Evaluate. 3/7x+6/x^2

OpenStudy (campbell_st):

needs a common denominator of 7x^2 \[\frac{3\times x}{7x \times x} + \frac{7\times6}{7\times x^2}= \frac{3x + 42}{7x^2}\]

OpenStudy (anonymous):

thanks